# Graphon Mixtures

### v1

@@ 1 Introduction
Observed sparse graphs such as social networks exhibit two contrasting types of behavior: a small number of large hubs inducing sparsity and a large number of small, dense communities (Zang et al., 2018; Zhao et al., 2021). Such hubs and communities are also observed in complex multi-functional networks such as neurological networks (Van den Heuvel & Sporns, 2013; Schwarz et al., 2008). A hub, viewed in isolation can be thought of as a star graph and a community by itself can be thought of as a dense graph. The prevalence of these two components in both biological and social networks is evidence that they are fundamental structures in evolving graphs.
@@
Motivated by these considerations we consider graph mixtures – a combination of two graphs, one sparse and one dense generated from two different types graph generators: graphons. A graphon is a symmetric and measurable function $W:[0,1]^{2}\rightarrow[0,1]$. Graphons are used to learn the underlying structure or representation of graphs (Xu et al., 2021) and recently graphons are used in neural operators and graph neural networks (Levie, 2023; Tieu et al., 2024; Cheng & Peng, 2023). A graphon can be thought of as a graph blueprint and can be used to generate graphs of an arbitrarily large size. As a consequence of the Aldous-Hoover theorem (Aldous, 1981), graphs generated from a graphon $W$ are dense. Thus, while it straightforward to generate dense graphs from a graphon $W$, generating sparse graphs from a graphon is not quite that simple.
@@
There have been several extensions to graphons to model sparse graphs. Caron et al. (2023) model graphs as exchangeable point processes and extend the classical graphon framework to the sparse regime. Their graphons are defined on $\mathbb{R}_{+}^{2}$ instead of on the unit square $[0,1]^{2}$. Research led by Borgs and Chayes tackle sparsity in different ways. For example, Borgs et al. (2018) consider ‘stretched’ and ‘rescaled’ graphons that can represent sparse graphs. Veitch & Roy (2015) introduce graphexes, a triple describing isolated edges, infinite stars and a graphon defined on $\mathbb{R}_{+}^{2}$. Kandanaarachchi & Ong (2024) model sparse graphs by considering graphons of line graphs. We use results from Kandanaarachchi & Ong (2024) and Janson (2016) to define graphon mixtures, which can generate sparse and dense graphs depending on the mixture properties.
@@
Our approach for sparse graph generation considers line graphs. Line graphs, also known as non-backtracking graphs (Krzakala et al., 2013) are obtained by mapping edges to vertices. A line graph of a graph $G_{n}$ denoted by $L(G_{n})$ maps edges of $G_{n}$ to vertices of $L(G_{n})$ and connects vertices in $L(G_{n})$ if the corresponding edges in $G_{n}$ share a vertex (Figure 1a). Kandanaarachchi & Ong (2024) showed that for a subset of sparse graphs, which they called square-degree, the line graphs are dense. Hence, for square-degree graphs, by taking line graphs, we are back in the space of dense graphs which we know how to model using the standard graphon. However, this hardly tells us anything about the structure of the line graph graphon. Janson (2016) showed that line graph limits are disjoint clique graphs and further showed that disjoint clique graphs can be modelled by a sequence of positive numbers, which he called a mass-partition. This result combined with the inverse line graph operation enable us to define graphon mixtures.
@@
The contributions of this paper are: (1) we propose graphon mixtures – a novel approach that explicitly models sparsity and thereby generates graph mixtures; (2) given a graph mixture, we estimate the degree of hubs of unseen graphs and the graphon corresponding to the sparse components, and we provide theoretical guarantees on our estimates that have exponential or polynomial convergence; (3) we show empirically that our approach is useful on synthetic and real data.
@@ 2 Limits of sparse graphs and line graphs
A graphon is a symmetric and measurable function $W:[0,1]^{2}\rightarrow[0,1]$. Graphs and graphons are linked by the adjacency matrix. The empirical graphon (Definition A.2) of a graph is obtained by replacing the adjacency matrix with a unit square where the $(i,j)$th entry of the adjacency matrix is replaced with a square of size $1/n\times 1/n$. Given a graphon $W$, a graph with $n$ nodes for any $n\in{\mathbb{N}}$ can be generated as follows:
@@ 2.1 Sparse graphs are stars in the limit
To address this problem, more mathematically intricate constructions and convergences have been defined (Borgs et al., 2018; Caron et al., 2023). Kandanaarachchi & Ong (2024) showed that if sparse graphs satisfy a certain condition, then their line graphs (Definition A.1, Figure 1a) are dense. They called this condition the square-degree property (Definition A.9). The set of graph sequences satisfying the square-degree property is denoted by $S_{q}$. They showed that if $\{G_{n}\}_{n}\in S_{q}$ and the line graphs $\{L(G_{n})\}_{n}$ converge, then they converge to a graphon $U\neq 0$. Here, we elucidate the structure of $U$, based on an earlier result by Janson (2016).
@@
A clique is a subset of nodes in the graph that are connected to each other. This result is important because as shown in Figure 1b, the inverse line graph of a disjoint clique graph is a disjoint star graph, and stars are the most sparse of the graphs that exhibit power-law degrees. Therefore, the line graph limit explains sparsity in a fundamental way via the inverse line graph operation. It is worth noting that if a graph is a line graph, then its inverse exists and is unique (Whitney, 1932) with the only exception when the line graph is a triangle. Janson (2016) represented disjoint clique graph limits by a sequence of proportions called the mass-partition (Definition A.5) where the $j$th element of the mass-partition gives the proportion of nodes in the $j$th disjoint clique. A mass-partition $\bm{p}=\{p_{i}\}_{i=1}^{\infty}$ uniquely defines a disjoint clique graphon (Definition A.6). We combine disjoint clique graph limits of (Janson, 2016) with converging sparse square-degree (Definition A.9) graphs of (Kandanaarachchi & Ong, 2024) in the following lemma. Furthermore, we provide a new definition (so called max-degree) which is equivalent to square-degree property.
@@ 2.2 Related work
Sparse graphs including scale-free and power-law graphs have been studied in depth. For scale-free graphs described by the Barabási-Albert model (Barabási & Albert, 1999) the proportion $P(d)$ of vertices with degree $d$ obeys a power law $P(d)\propto d^{-\gamma}$, where $\gamma\approx 3$. Bollobás et al. (2001) showed for these graphs the maximum degree is $\Theta(\sqrt{n})$, where $n$ is the number of the nodes in the graph. For these graphs the probability $\Pi(i)$ that a new node connects to node $i$, which has degree $d_{i}$ is given by $\Pi(i)=\frac{d_{i}^{\alpha}}{\sum_{d_{i}^{\alpha}}}$ where $\alpha=1$. When $\alpha>1$, the graphs are described by a different regime called the super-linear preferential attachment. For these graphs, the maximum degree $d_{\max}\to n$ exhibiting a winner takes it all strategy while other degrees remain bounded (Sethuraman & Venkataramani, 2019). Thus, for $\alpha=1$ the maximum degree $d_{\max}\in\Theta(\sqrt{n})$ and for $\alpha>1$ , $d_{\max}\to n$.
In contrast, classical graphons were used to describe dense graphs. Over time the original definitions, constructions and convergences were extended to incorporate sparse graphs (Borgs et al., 2021; Caron et al., 2023). One such construction is graphexes. Graphexes describe a triple $(I,S,W)$, where $I$ describes isolated edges, $S$ infinite stars and $W$ a graphon defined on $\mathbb{R}_{+}^{2}$. This particular graphon $W$ can generate dense or sparse graphs. For dense graphs, the graphon is defined on a compact $S\subset\mathbb{R}_{+}^{2}$, while for sparse graphs $W$ is defined on $\mathbb{R}_{+}^{2}$. Thus, extending the support of the graphon from a finite subset to the whole positive quadrant enabled sparsity in this instance.
Our approach is different as we use inverse line graphs of disjoint clique graphs to generate the sparse component. As both graphons in our mixture are defined on the unit square, it makes it easier to estimate these graphons in practice. Furthermore, graphon mixtures give rise to top-$k$ degrees in between the Barabási-Albert model and the super-linear preferential attachment graphs.
@@ 3 Mixtures of sparse and dense graphs
We generate a graph mixture first by picking a graphon $W$ and a disjoint clique graphon $U$. Graphon $U$ lives in the line graph space and is the key to generating sparse graphs. By taking the inverse line graph of graphs generated by $U$, we get sparse graphs. Thus, $U$ models sparsity by the inverse line graph operation.
First we generate two graph sequences using $U$ and $W$, noting that both graph sequences are dense, but the graph sequence generated by $U$ lives in the line graph space. The sparse graph sequence is obtained by taking the inverse line graphs of the graphs generated by $U$. Then we join the $i$th graph generated by $W$ with the $i$th inverse line graph generated by $U$ according to a set of simple rules and obtain the graph mixture. A key component in the mixing process is the ratio of the number of nodes in the sparse graph to that of the dense graph. If this ratio goes to infinity along the sequence, then the mixture sequence is sparse. If the ratio is bounded, then the mixture sequence is dense. Thus, for a given $W$ and $U$ by changing the mixture ratio function, we can generate graph sequences ranging from dense to sparse.
@@ 3.1 The graph mixture model
Given two graph sequences, we define joining rules that are not too restrictive.
Definition 3.1.
  1. 1.
    No new nodes condition: no new nodes are added as part of the joining process.
  2. 2.
    No deletion condition: No nodes or edges are deleted as part of the joining process.
  3. 3.
    Random edges condition: if edges are added, they are added randomly in the sense that nodes within a graph are equally likely to be selected to form edges.
  4. 4.
    New edges condition: if edges are added, the number of new edges added $m_{new_{i}}$ satisfies $m_{new_{i}}=cm_{d_{i}}$, for some $c\in\mathbb{R}$, where $m_{d_{i}}$ denotes the number of edges in $G_{d_{i}}$.
The graph joining rules ensure that each node in $G_{n_{i}}$ comes either from $G_{s_{i}}$ or $G_{d_{i}}$. Furthermore, the randomness in the joining process prevents creating additional structure in the graph.
@@ 3.2 Expectations and edge density
Lemma 3.4 shows that $(U,W)$-mixture graphs are a general construction that allows us to generate sparse or dense graphs. While our focus is on sparse graphs, dense graphs that cannot be fully explained by a single graphon $W$ can be generated by a $(U,W)$ mixture. Thus, $(U,W)$-mixture graphs can model a richer set of dense graphs compared to a single graphon $W$. We briefly discuss this in Appendix B and give examples of $(U,W)$-mixture graphs that range from dense to sparse using the same $(U,W)$ combination.
@@ 4 The top-$k$ degrees of sparse $(U,W)$ mixture graphs
Suppose $G_{n_{i}}$ and $G_{n_{j}}$ are two graphs from a sparse $(U,W)$-mixture graph sequence (Definition 3.3). We treat $G_{n_{i}}$ as the training graph and $G_{n_{j}}$ as the test graph. Suppose $G_{n_{i}}$ and $G_{n_{j}}$ have $n_{i}$ and $n_{j}$ nodes respectively. Let $\bm{p}=(p_{1},p_{2},\ldots)$ be the mass-partition (Definition 3.2) associated with $U$, which has at least $k$ partitions. Let $\text{deg }_{G_{n_{i}}}v_{(\ell)}$ denote the $\ell$th largest degree in $G_{n_{i}}$ where $\ell\leq k$. Then we estimate the $\ell$th largest degree in $G_{n_{j}}$ as
@@ 5 Estimating $U$ for finite number of partitions
Understanding sources of sparsity can help uncover the graph generation mechanism. For mixture graphs, sparsity comes from $U$, which lives in the line graph space and can have finite or infinite partitions (Definition 3.2). By observing mixture graphs, we estimate $\bm{p}$, both when $\bm{p}$ has finite and infinite partitions. In this section we consider finite partitions. For both cases, we consider a class of well-behaved $W$ in our mixture model. We do not need $W$ to be well-behaved to estimate the topmost degrees of $G_{n_{i}}$. However, we need some conditions on $W$ so that we can estimate $U$.
@@
When $W$ satisfies Assumption 5.2, the degrees in the dense part $G_{d_{i}}$ are closely packed, i.e., given a degree value, there is another degree value quite close in $G_{d_{i}}$. We make this notion precise in Lemma D.4, which we use to estimate $k$, the number of partitions.
@@
Let $\{G_{n_{i}}\}_{i}$ be a sequence of sparse $(U,W)$-mixture graphs (Definition 3.3) with dense and sparse parts $G_{d_{i}}$ and $G_{s_{i}}$ having nodes $n_{d_{i}}$ and $n_{s_{i}}$ respectively. Suppose $W$ satisfies Assumption 5.2 and $W\neq 1$. Let $\bm{p}=(p_{1},p_{2},\ldots)$ be the mass-partition (Definition 3.2) associated with $U$ which has only $k$ partitions. Let $\text{deg }_{G_{n_{i}}}v_{(\ell)}$ denote the $\ell$th largest degree in $G_{n_{i}}$. Then there exists $I_{0}$ such that for $i>I_{0}$ we have
@@
where $\text{deg }_{G_{n_{i}}}v_{(\ell)}$ denotes the observed $\ell$th highest degree in ${G_{n_{i}}}$. From the Order Preserving Property (Proposition 4.1) we know that for large $i$, $\text{deg }_{G_{n_{i}}}v_{(\ell)}=q_{\ell,i}$, and thus the bounds in Proposition 5.4 are satisfied.

### v2

@@ 1 Introduction
Observed sparse graphs such as social networks exhibit two contrasting types of behavior: a small number of high degree nodes called hubs inducing sparsity and a large number of small, dense communities (Zang et al., 2018; Zhao et al., 2021). Such hubs and communities are also observed in complex multi-functional networks such as neurological networks (Van den Heuvel & Sporns, 2013; Schwarz et al., 2008). A hub, viewed in isolation can be thought of as a star graph and a community by itself can be thought of as a dense graph. The prevalence of these two components in both biological and social networks is evidence that they are fundamental structures in evolving graphs.
@@
Motivated by these considerations we consider graph mixtures – a combination of two graphs, one sparse and one dense generated from two different types of graph generators: graphons. A graphon is a symmetric and measurable function $W:[0,1]^{2}\rightarrow[0,1]$. Graphons are used to learn the underlying structure or representation of graphs (Xu et al., 2021) and recently graphons are used in neural operators and graph neural networks (Levie, 2023; Tieu et al., 2024; Cheng & Peng, 2023). A graphon can be thought of as a graph blueprint and can be used to generate graphs of an arbitrarily large size. As a consequence of the Aldous-Hoover theorem (Aldous, 1981), graphs generated from a graphon $W$ are dense. Thus, while it straightforward to generate dense graphs from a graphon $W$, generating sparse graphs from a graphon is not quite that simple.
@@
There have been several extensions to graphons to model sparse graphs. Caron et al. (2023) model graphs as exchangeable point processes and extend the classical graphon framework to the sparse regime. Their graphons are defined on $\mathbb{R}_{+}^{2}$ instead of on the unit square $0,1]^{2}$. Research led by Borgs and Chayes tackle sparsity in different ways. For example, Borgs et al. (2018) consider ‘stretched’ and ‘rescaled’ graphons that can represent sparse graphs. Veitch & Roy (2015) introduce graphexes, a triple describing isolated edges, infinite stars and a graphon defined on $\mathbb{R}_{+}^{2}$. Kandanaarachchi & Ong (2024) model sparse graphs by considering graphons of line graphs. We use results from Kandanaarachchi & Ong (2024) and Janson (2016) to define graphon mixtures, which can generate sparse and dense graphs depending on the mixture properties.
@@
Our approach for sparse graph generation considers line graphs. Line graphs, also known as non-backtracking graphs (Krzakala et al., 2013) are obtained by mapping edges to vertices. A line graph of a graph $G_{n}$ denoted by $L(G_{n})$ maps edges of $G_{n}$ to vertices of $L(G_{n})$ and connects vertices in $L(G_{n})$ if the corresponding edges in $G_{n}$ share a vertex (Figure 1(a)). Kandanaarachchi & Ong (2024) showed that for a subset of sparse graphs, which they called square-degree, the line graphs are dense. Hence, for square-degree graphs, by taking line graphs, we are back in the space of dense graphs which we know how to model using the standard graphon. However, this hardly tells us anything about the structure of the line graph graphon. Janson (2016) showed that line graph limits are disjoint clique graphs and further showed that disjoint clique graphs can be modelled by a sequence of positive numbers, which he called a mass-partition. This result combined with the inverse line graph operation enable us to define graphon mixtures.
@@
The contributions of this paper are: (1) we propose graphon mixtures – a novel approach that explicitly models sparsity and thereby generates graph mixtures; (2) given a graph mixture, we estimate the degree of hubs of unseen graphs and the graphon corresponding to the sparse components, and we provide theoretical guarantees on our estimates that have exponential or polynomial convergence; (3) we show empirically that our approach is useful on synthetic and real data. All proofs are presented in the Appendix.
@@ 2 Limits of sparse graphs and line graphs
A graphon is a symmetric and measurable function $W:[0,1]^{2}\rightarrow[0,1]$. Graphs and graphons are linked by the adjacency matrix. The empirical graphon (Definition A.2) of a graph is obtained by scaling the adjacency matrix to the unit square and coloring smaller squares of size $1/n\times 1/n$ black if the corresponding entry in the adjacency matrix is 1 and coloring it white if the entry is 0. Given a graphon $W$, a graph with $n$ nodes for any $n\in{\mathbb{N}}$ can be generated as follows:
@@ 2.1 Sparse graphs are stars in the limit
To address this problem, more mathematically intricate constructions and convergences have been defined (Borgs et al., 2018; Caron et al., 2023). Kandanaarachchi & Ong (2024) showed that if sparse graphs satisfy a certain condition, then their line graphs (Definition A.1, Figure 1(a)) are dense. They called this condition the square-degree property (Definition A.9). The set of graph sequences satisfying the square-degree property is denoted by $S_{q}$. They showed that if $\{G_{n}\}_{n}\in S_{q}$ and the line graphs $\{L(G_{n})\}_{n}$ converge, then they converge to a graphon $U\neq 0$. Here, we elucidate the structure of $U$, based on an earlier result by Janson (2016).
@@
A clique is a subset of nodes in the graph that are connected to each other. This result is important because as shown in Figure 1(b), the inverse line graph of a disjoint clique graph is a disjoint star graph, and stars are the most sparse of the graphs that exhibit power-law degrees. Therefore, the line graph limit explains sparsity in a fundamental way via the inverse line graph operation. It is worth noting that if a graph is a line graph, then its inverse exists and is unique (Whitney, 1932) with the only exception when the line graph is a triangle. Janson (2016) represented disjoint clique graph limits by a sequence of proportions called the mass-partition (Definition A.5) where the $j$th element of the mass-partition gives the proportion of nodes in the $j$th disjoint clique. A mass-partition $\bm{p}=\{p_{i}\}_{i=1}^{\infty}$ uniquely defines a disjoint clique graphon $W_{\bm{p}}^{\mathcal{M}}$ (Definition A.6) resembling a block-diagonal matrix (see $U$ in Figure 2). We combine disjoint clique graph limits of Janson (2016) with converging sparse square-degree graphs $S_{q}$ (Definition A.9) of Kandanaarachchi & Ong (2024) in the following lemma. Furthermore, we provide a new definition (so called max-degree) which is equivalent to square-degree property.
@@ 2.2 Related work
Sparse graphs including scale-free and power-law graphs have been studied in depth. For scale-free graphs described by the Barabási-Albert model (Barabási & Albert, 1999) the proportion $P(d)$ of vertices with degree $d$ obeys a power law $P(d)\propto d^{-\gamma}$, where $\gamma\approx 3$. Bollobás et al. (2001) showed for these graphs the maximum degree is $\Theta(\sqrt{n})$, where $n$ is the number of the nodes in the graph.
Leskovec et al. (2010) proposed a graph generative model called Kronecker graphs. They modeled graph evolution using the Kronecker product and had two versions (deterministic and stochastic) of graph generators. They showed they can generate graphs with heavy-tailed degree distributions and other such properties.
Classical graphons were used to describe dense graphs. Over time the original definitions, constructions and convergences were extended to incorporate sparse graphs (Borgs et al., 2021). Caron & Fox (2017) and Caron et al. (2023) modeled graphs as exchangeable point processes where the graphon $W$ is defined on $\mathbb{R}_{+}^{2}$. These were also known as graphex processes. They used Kallenberg exchangeability in their theoretical framework and sparsity was achieved by extending the support of the graphon from a finite subset to the whole positive quadrant.
Our approach is different to the above as we use inverse line graphs of disjoint clique graphs to generate the sparse component. As both graphons in our mixture are defined on the unit square, it makes it easier to estimate these graphons in practice, especially the sparse component. By varying mixture properties we can generate both dense and sparse graphs.
@@ 3 Mixtures of sparse and dense graphs
Figure 2: The overview of $(U,W)$-mixture graphs. A disjoint clique graph $H_{m}$ is sampled from graphon $U$. The inverse line graph $G_{s}=L^{-1}(H_{m})$ is a disjoint set of stars. A graph $G_{d}$ is sampled from graphon $W$. Then, $G_{s}$ and $G_{d}$ are joined according to graph joining rules resulting in the mixture graph $G_{n}$.
We generate a graph mixture by picking a graphon $W$ and a disjoint clique graphon $U$ (see Figure 2). First we generate two graph sequences using $U$ and $W$, noting that both graph sequences are dense, but the graph sequence generated by $U$ lives in the line graph space. The sparse graph sequence is obtained by taking the inverse line graphs of the graphs generated by $U$. Then we join the $i$th graph generated by $W$ with the $i$th inverse line graph generated by $U$ according to a set of simple joining rules and obtain the graph mixture. The results in Sections 4, 5 and 6 do not have any terms pertaining to the join. They hold if we consider the disjoint union of the two graph sequences. However, the real graphs in Table 2 are not disjoint unions. As such, we consider a simple way to join nodes randomly without adding a lot of extra edges. We ensure the main structural parts are contributed by $U$ and $W$, but not the join.
A key component in the mixing process is the ratio of the number of nodes in the sparse graph to that of the dense graph. If this ratio goes to infinity along the sequence, then the mixture sequence is sparse. If the ratio is bounded, then the mixture sequence is dense. Thus, for a given $W$ and $U$ by changing the mixture ratio function, we can generate graph sequences ranging from dense to sparse.
@@ 3.1 The graph mixture model
Given two graph sequences – one dense and one sparse – we consider a simple way to randomly join a dense graph and a sparse graph with a given number of edges, making sure the main structural parts are contributed by the two graphs but not the join. We formally define the joining rules below.
Definition 3.3.
  1. 1.
    No new nodes: no new nodes are added as part of the joining process.
  2. 2.
    No deletion: No nodes or edges are deleted as part of the joining process.
  3. 3.
    Random edges: if edges are added, they are added randomly in the sense that nodes within a graph are equally likely to be selected to form edges.
  4. 4.
    New edges: if edges are added, the number of new edges added $m_{new_{i}}$ satisfies $m_{new_{i}}=cm_{d_{i}}$, for small $c\in\mathbb{R}$, where $m_{d_{i}}$ denotes the number of edges in $G_{d_{i}}$.
In defining these rules our motivation is to keep the joining process as simple as possible. As such, we have employed a random join. These rules can be further explored and modified to suit different needs that result in more targeted mixtures.
@@ 3.2 Expectations and edge density
Lemma 3.4 shows that $(U,W)$-mixture graphs are a general construction that allows us to generate sparse or dense graphs. While our focus is on sparse graphs, dense graphs that cannot be fully explained by a single graphon $W$ can be generated by a $(U,W)$ mixture (see Appendix B.2). Thus, $(U,W)$-mixture graphs can model a richer set of dense graphs compared to a single graphon $W$. We give examples of $(U,W)$-mixture graphs in Appendix B and show that by changing mixture properties we get graph sequences ranging from dense to sparse using the same $(U,W)$ combination.
@@ 4 The top-$k$ degrees of sparse $(U,W)$ mixture graphs
[…]
@@ 5 Estimating $U$ for finite number of partitions
Graphon $U$ describes the relative strength (or vulnerability) of the hubs in a mixture. In a social media network, hubs act as superspreaders of information and reach; $U$ shows the relative power of the big players in the network. In an energy/telecommunications network, it can show the vulnerability/resilience in the event of an attack. Consider two energy networks $A$ and $B$ with $U_{A}=(0.5,0.3,0.2)$ and $U_{B}=(0.1,\ldots,0.1)$. Then mixture $A$ is far more vulnerable from a resilience viewpoint compared to mixture $B$ if attacked. By attacking one node (the hub corresponding to 0.5), approximately 50% of the network can be crippled in an energy/telecommunications network scenario.
Graphon $U$ can have finite or infinite partitions. By observing mixture graphs, we estimate the mass-partition $\bm{p}$ (Definition 3.1), both when $\bm{p}$ has finite and infinite partitions. In this section we consider finite partitions. For both cases, we consider a class of well-behaved $W$ in our mixture model. We do not need $W$ to be well-behaved to estimate the topmost degrees of $G_{n_{i}}$. However, we need some conditions on $W$ so that we can estimate $U$.
@@
When $W$ satisfies Assumption 5.2, the degrees in the dense part $G_{d_{i}}$ are closely packed, i.e., given a degree value, there is another degree value quite close in $G_{d_{i}}$ (see Lemma D.4). Using Assumption 5.2 we estimate $k$, the number of partitions.
@@
[…]
@@
[…]